# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 5.4 Extragalactic contamination
We quantify the extragalactic contamination present in our dust models using a tomographic redshift-clustering technique (Schmidt et al., 2015; Chiang & Ménard, 2019). Our intensity-based Galactic dust templates inevitably contain emission from both Galactic dust and external galaxies. As described in Section 3.1.3, the new d9 and d10 dust templates are derived from GNILC-processed Planck data, while older PySM dust templates used Commander data products. We thus expect that the new Galactic dust models are significantly less affected by CIB contamination than previous models. Here we quantify this contamination by measuring the cross-correlation between our dust models and the clustering of galaxies as a function of redshift in spectroscopic survey data. A perfect Galactic dust template would be uncorrelated with such clustering; the signature of CIB contamination is excess template emission correlated with galaxy clustering.
@@
Following a procedure similar to Chiang & Ménard (2019), we compute the cross-correlation between local fluctuations in the Galactic emission maps and galaxy density maps. The latter are constructed by stacking the number of galaxies from the GLADE+ catalog (Dálya et al., 2022) in $N_{\rm side}=2048$ HEALPix pixels within redshift bins of $\Delta z\sim 0.03$ over the range $0.350<z<$ 0.35, and then smoothing the resulting maps to $22\arcmin$. We compute this cross-correlation for each of the d1, d10, and d12 dust emission templates at 857 GHz, at high Galactic latitudes (in the GAL70 mask) within the catalog footprint. Figure 12 shows that while each of the new GNILC-based maps contain less extragalactic contamination than d1, the decreased contamination is more marked in d10 than in d12. This is expected, as d10 and similar templates have random small-scale structure, while d12 retains more of the structure of the data at small scales. Although the shape of the correlation curves is affected by the galaxy distribution and catalog completeness as a function of redshift, the GLADE+ catalog is up to 90% complete at $\sim 500~{}\mathrm{Mpc}$ ($z\sim 0.1$; Dálya et al., 2022).
@@
We perform the same cross-correlation analysis to quantify the extragalactic contamination in the $\beta_{d}$ maps. We find, as expected, higher extragalactic contamination in the d1 $\beta_{d}$ map than in the d10 $\beta_{d}$ map. We emphasize that the mitigation of CIB contamination in the GNILC-processed Planck data affects both the spatial structure of the template frequency maps and their extrapolation to other frequencies via the spectral parameter maps.
5.5 Non-Gaussianity
@@
Figure 13: MFs for the small scales of three sets of $Q$ maps on the sphere with GAL080 mask. The large scales are filtered out by excluding multipoles $\ell<200$ in the maps, and we choose $\ell_{\text{max}}=2048$ when obtaining the small scales. We show the MFs as a function of threshold $\rho$, for Gaussian-mod-ss in blue, poltens-ss in orange and Gaussian-ss in dashed gray.
@@
We quantify the level of non-Gaussianity in the small-scale dust emission generated through the polarization fraction tensor transformation used in dust model d10 at 353 GHz (which is identical to d9 at that frequency). To measure the non-Gaussianity in the maps we consider the Minkowski Functionals (MFs, Minkowski, 1903), which are a common tool to quantify map-space, higher-order statistical properties (e.g., Rahman et al., 2021; Martire et al., 2023; Carones et al., 2024). Hadwiger’s theorem implies that, for any $n$-dimensional excursion set defined with a threshold value $\rho$, there exist $n+1$ MFs that geometrically and topologically describe the morphology of the set (Hadwiger, 1957). In our case, for two-dimensional maps, we have three kinds of MFs. These are $\mathcal{V}_{0}$, $\mathcal{V}_{1}$, and $\mathcal{V}_{2}$, which correspond to the area, the perimeter and the connectivity of the excursion set (iso-intensity contour), respectively.
@@
We use these MFs to compare the small-scale structure in d10 to those of two other sets of maps: (i) maps where the small-scale structures are fully Gaussian and isotropic and (ii) maps where the small-scale structures are Gaussian but anisotropic across the sky. As described Section 3.1.2, the small scale structures in the d10 model are generated as a Gaussian random field in $i$, $q$, and $u$, and are multiplied by the $m_{i}$ and $m_{p}$ modulation maps before they are coadded to the large-scale maps and transformed back into $I$, $Q$ and $U$. We want to understand the impact of this effective modulation on the MFs, and therefore construct a set of maps that are amplitude-modulated versions of Gaussian-random-field maps. This allows us to disentangle any non-Gaussianity generated through the modulation from the potential non-Gaussianity introduced due to the polarization fraction tensor transformation.
@@
The first set of maps contains isotropic small-scale structure. We construct the small-scale structure by generating a Gaussian random field with power-law power spectra in $TT$, $EE$, and $BB$, using power laws fit to the power spectrum of the d10 maps on the GAL097 mask in the multipole range $[800,2000]$. This Gaussian random field is then high-pass filtered to remove power below $\ell_{cut}=200$ in both total intensity and polarization, and co-added to the large-scale dust template.
@@
For the second set of maps, we introduce a modulation in $I$, $Q$, and $U$ that mimics the effective modulation applied to $i$, $q$, and $u$ in the construction of the d10 maps. We generate modulation maps $m_{I}$ and $m_{P}$ from $IQU$, following the same procedure as done in $iqu$ space (Equations 89). We multiply the Gaussian isotropic small-scale map described above with these modulation maps, and then co-add them with the large-scale dust template. However, we want to ensure that the power spectra of modulated Gaussian small scales computed on the individual sky-fraction masks are as close as possible to the d10 map. To achieve this, we adjust the modulation maps $m_{I}$ and $m_{P}$ by applying different multiplicative factors to non-overlapping regions of the sky. We multiply $m_{I}$ and $m_{P}$ on the GAL40 mask by scalar factors that adjust the mean power spectrum of the modulated small-scale structure to be equal to the power spectrum of d10 on the same mask. We repeat this process for each successively larger sky-fraction mask, applying the approximately order-unity multiplication factor to the non-overlapping sky region at each iteration.
@@
We thus consider three sets of maps with different small scales co-added: model d10, a map with purely Gaussian small-scale structure, and a map with modulated Gaussian small-scale structure. We refer to these three maps as poltens-ss, Gaussian-ss, and Gaussian-mod-ss, respectively. We apply a high-pass filter with $\ell_{min}=200$, using a smooth function similar to Equation 11, to retain only the small scales of these maps. We calculate the MFs both on the sphere and in several selected regions of the sky projected into a Cartesian projection.
5.5.1 Minkowski Functionals on the sphere
@@
Following the algorithm in Grewal et al. (2022), we calculate the MFs for the three $Q$ maps on the sphere, i.e., in HEALPix format, on the GAL080 mask. We first normalize the maps by dividing each map by its standard deviation, and compute the MFs for iso-intensity contours in the range $[-3,3]$ (Figure 13). The MFs of the corresponding $U$ maps look very similar to $Q$ maps and are not shown here.
@@
Figure 13 shows that when averaging over a large sky area, the MFs of Gaussian-mod-ss and poltens-ss are almost identical, while the MFs of Gaussian-ss differ substantially. The difference in MFs between poltens-ss and Gaussian-ss indicates the existence of non-Gaussianity in poltens-ss, but the similarity between poltens-ss and Gaussian-mod-ss demonstrates that the non-Gaussianity in poltens-ss comes from the anisotropy in the maps, which originates from the modulation, rather than from the polarization fraction tensor transformation.
@@
Figure 14: Zoom-in plots of the selected patch centered at $(l,b)=(-15^{\circ}$, $45^{\circ})$. From left to right: the final map with Gaussian-mod-ss, final map with poltens-ss (i.e., d10 map), Gaussian-mod-ss only map and poltens-ss only map. From top to bottom we show $I$, $Q$, and $U$, respectively. The colorbar on the left indicates the pixel values in the left-most two columns in the units of $\mu\rm K_{RJ}$ and the colorbar on the right is for the last two columns.
@@
Figure 15: Minkowski Functionals as a function of the threshold $\rho$ for one realization of $I$, $Q$ and $U$ small scales in the patch centered on $(l,b)=(-15^{\circ}$, $45^{\circ})$ in Galactic coordinates. Each row shows three kinds of Minkowski Functionals. The blue dotted one is for Gaussian-mod-ss while the orange solid one is from poltens-ss. We also show the Gaussian-ss in dashed gray line as a comparison.
5.5.2 Minkowski Functionals on small regions
@@
We consider a $20^{\circ}\times 20^{\circ}$ region centered at $(l,b)=(-15^{\circ}$, $45^{\circ})$ to determine whether significant differences in the MFs between Gaussian-mod-ss and poltens-ss sets of maps exist in small regions of sky. Those maps are shown in Figure 14. We can see by eye that poltens-ss contains structure that is not present in the Gaussian-mod-ss maps. We calculate the MFs of these small-scale maps, following Mantz et al. (2008) for the calculation of MFs for a square patch. Before the calculation, we also rescale all the small scales linearly to be in the range $[-1,1]$.
@@
Figure 15 shows the MFs of the Gaussian-mod-ss and poltens-ss maps presented in Figure 14. In contrast with the large-area results presented in Figure 13, in this case we do measure a departure of the poltens-ss MFs from the Gaussian-mod-ss ones. This means that the polarization fraction tensor transformation introduced non-Gaussian small-scale structure, distinct from pure modulation effects, that is detectable over small sky regions. We conclude that the level of induced non-Gaussianity differs from region to region and does not significantly impact the statistical properties when averaged over large sky fractions.
6 Future Outlook
@@
The Galactic emission templates employed in this work are derived from component separation of the microwave sky and thus are subject to both noise and fitting degeneracies. This is particularly evident in the spectral parameter maps. The $T_{d}$ and $\beta_{d}$ template maps have a strong anti-correlation that can arise from fitting degeneracies, particularly at low signal-to-noise ratios (Shetty et al., 2009). These maps also exhibit correlations with the dust intensity maps, particularly in diffuse, low-signal regions. While all dust models developed in this work respect current observational constraints on frequency decorrelation (see Section 5.3), it is likely that some variation in the spectral parameter maps is attributable to noise rather than true astrophysical variation. Construction of improved spectral parameter maps is feasible by employing improved algorithms on existing data (e.g., Watts et al., 2024) and should be a high priority for future work.
@@
By imposing a uniform cut-off in $\ell$ over which data-driven templates are employed, the models developed here are discarding high signal-to-noise information at small scales over portions of the sky where emission is bright. Future methodologies would ideally retain information from the templates on as small a scale as possible, with the cut-off scale varying across the sky. Artifacts from the splicing of large-scale templates with small-scale synthetic structures, including the ratio of spatial power across the modulation scale discussed in Section 5.2.2, should be further mitigated. We note that the current methodology is fundamentally not designed to be able to reproduce the observed power spectra of all quantities over all sky areas, but future approaches could improve the behavior of the power spectrum beyond the modulation scale and for smaller sky masks. Further, particularly as data are being collected over partial sky areas by sensitive ground-based experiments, methods to incorporate partial sky templates should be developed.
@@
The current models do not explicitly impose any non-zero parity-odd correlations in the polarized dust emission, i.e., the $TB$ and $EB$ correlations are zero in the power spectra that are used to extrapolate the small-scale dust emission structure. However, analysis of the Planck polarized dust emission finds a significant nonzero $TB$ correlation at $\ell\lesssim 600$ (Planck Collaboration XI, 2020; Weiland et al., 2020). A proposed physical mechanism for the origin of the nonzero $TB$ signal is misalignment between dusty filaments and the projected magnetic field orientation (Huffenberger et al., 2020; Clark et al., 2021; Cukierman et al., 2023): this picture also predicts the sign and amplitude of an expected nonzero $EB$ correlation. Future work could incorporate this parity-odd contribution to the polarized dust emission (e.g., Hervías-Caimapo et al., 2024). Such models would be of particular use for the development of analysis techniques that seek to constrain signatures of cosmic birefringence or other parity-violating physics in the CMB (e.g., Choi et al., 2020; Minami & Komatsu, 2020; Eskilt & Komatsu, 2022; BICEP/Keck Collaboration et al., 2024).
@@
Future work could also improve the physical realism of the small-scale emission structure. The structure of polarized dust emission on small scales is highly filamentary (e.g., Clark et al., 2015; Halal et al., 2024a). Realizations of small-scale structure that have the particular character of non-Gaussianity that is realized in the sky could be generated using, for example, generative adversarial neural networks (Krachmalnicoff & Puglisi, 2021; Yao et al., 2024) or scattering-transform-based techniques (Regaldo-Saint Blancard et al., 2020; Delouis et al., 2022; Mousset et al., 2024), which could be informed by the non-Gaussian statistics of diffuse ISM gas tracers (Lei & Clark, 2023). Another approach is to explicitly model the observed emission as arising from a superposition of many interstellar filaments, perhaps informed by the structure of interstellar neutral hydrogen emission (Clark & Hensley, 2019), or as a flexible parametric model (Hervías-Caimapo & Huffenberger, 2022; Hervías-Caimapo et al., 2024).
@@
As our knowledge of diffuse ISM physics improves, there are further opportunities for more sophisticated modeling of the Galactic emission. A recently discovered link between the mass fraction of neutral hydrogen in the cold neutral medium and the 353 GHz dust polarization fraction implies that dust polarization models could be further improved by incorporating information on the phase structure of the neutral ISM (Lei & Clark, 2024). It was also recently found that the three-dimensional structure of nearby interstellar dust has a measurable imprint on the dust polarization structure at CMB-relevant frequencies (Halal et al., 2024b). Thus, as models of the 3D dust distribution (e.g., Edenhofer et al., 2024; Zhang & Green, 2024) improve, it may be possible to incorporate these data products into models of the dust polarization field. Starlight polarization data can further constrain the plane-of-sky magnetic field orientation of nearby dust structures (e.g., Pelgrims et al., 2024; Panopoulou et al., 2025).
@@
In addition to the small-scale structure of the diffuse dust and synchrotron emission, the real sky at these frequencies contains emission from discrete sources that are not explicitly included in our models. These include cold clumps (Clancy et al., 2023), as well as planetary nebulae, supernova remnants, pulsar wind nebulae, and so forth (Naess et al., 2020; Guan et al., 2021). Nearby galaxies, while not technically within the scope of Galactic emission modeling, should also be incorporated as they are not included in existing simulations of the extragalactic sky.
7 Recommended Model Suite

### v2

@@ V.4 Extragalactic contamination
We quantify the extragalactic contamination present in our dust models using a tomographic redshift-clustering technique (Schmidt:2015; Chiang:2019). Our intensity-based Galactic dust templates inevitably contain emission from both Galactic dust and external galaxies. As described in Section III.1.3, the new d9 and d10 dust templates are derived from GNILC-processed Planck data, while older PySM dust templates used Commander data products. We thus expect that the new Galactic dust models are significantly less affected by CIB contamination than previous models. Here we quantify this contamination by measuring the cross-correlation between our dust models and the clustering of galaxies as a function of redshift in spectroscopic survey data. A perfect Galactic dust template would be uncorrelated with such clustering; the signature of CIB contamination is excess template emission correlated with galaxy clustering.
@@
Following a procedure similar to Chiang:2019, we compute the cross-correlation between local fluctuations in the Galactic emission maps and galaxy density maps. The latter are constructed by stacking the number of galaxies from the GLADE+ catalog (Dalya:2022) in $N_{\rm side}=2048$ HEALPix pixels within redshift bins of $\Delta z\sim 0.03$ over the range $0<z<0.35$, and then smoothing the resulting maps to $22\arcmin$. We compute this cross-correlation for each of the d1, d10, and d12 dust emission templates at 857 GHz, at high Galactic latitudes (in the GAL70 mask) within the catalog footprint. Figure 12 shows that while each of the new GNILC-based maps contain less extragalactic contamination than d1, the decreased contamination is more marked in d10 than in d12. This is expected, as d10 and similar templates have random small-scale structure, while d12 retains more of the structure of the data at small scales. Although the shape of the correlation curves is affected by the galaxy distribution and catalog completeness as a function of redshift, the GLADE+ catalog is up to 90% complete at $\sim 500~\mathrm{Mpc}$ ($z\sim 0.1$; Dalya:2022).
@@
We perform the same cross-correlation analysis to quantify the extragalactic contamination in the $\beta_{d}$ maps. We find, as expected, higher extragalactic contamination in the d1 $\beta_{d}$ map than in the d10 $\beta_{d}$ map. We emphasize that the mitigation of CIB contamination in the GNILC-processed Planck data affects both the spatial structure of the template frequency maps and their extrapolation to other frequencies via the spectral parameter maps. CIB contamination in dust models can create spurious small-scale map structure, depress the map polarization fraction, and imply specious dust spectral parameters that affect the map frequency scaling.
V.5 Non-Gaussianity
@@
Figure 13: MFs for the small scales of three sets of $Q$ maps on the sphere with GAL080 mask. The large scales are filtered out by excluding multipoles $\ell<200$ in the maps, and we choose $\ell_{\text{max}}=2048$ when obtaining the small scales. We show the MFs as a function of threshold $\rho$, for a map employing unmodulated (Gaussian_ss, dashed gray) and modulated (Gaussian_mod_ss, dashed blue) Gaussian small scales. The poltens_ss (solid orange) MFs show the case for modulated non-Gaussian small scales obtained by high-pass filtering the d10 $Q$ map.
@@
We quantify the level of non-Gaussianity in the small-scale dust emission generated through the polarization fraction tensor transformation used in dust model d10 at 353 GHz (which is identical to d9 at that frequency). To measure the non-Gaussianity in the maps we consider the Minkowski Functionals (MFs, Minkowski1903), which are a common tool to quantify map-space, higher-order statistical properties (e.g., Rahman:2021; Martire:2023; Carones:2024). Hadwiger’s theorem implies that, for any $n$-dimensional excursion set defined with a threshold value $\rho$, there exist $n+1$ MFs that geometrically and topologically describe the morphology of the set (hadwigerVorlesungenUeberInhalt1957). In our case, for two-dimensional maps, we have three kinds of MFs. These are $\mathcal{V}_{0}$, $\mathcal{V}_{1}$, and $\mathcal{V}_{2}$, which correspond to the area, the perimeter and the connectivity of the excursion set (iso-intensity contour), respectively.
@@
We use these MFs to compare the small-scale structure in d10 to those of two other sets of maps: (i) maps where the small-scale structures are fully Gaussian and isotropic and (ii) maps where the small-scale structures are Gaussian but anisotropic across the sky. As described Section III.1.2, the small scale structures in the d10 model are generated as a Gaussian random field in $i$, $q$, and $u$, and are multiplied by the $m_{i}$ and $m_{p}$ modulation maps before they are coadded to the large-scale maps and transformed back into $I$, $Q$ and $U$. We want to understand the impact of this effective modulation on the MFs, and therefore construct a set of maps that are amplitude-modulated versions of Gaussian-random-field maps. This allows us to disentangle any non-Gaussianity generated through the modulation from the potential non-Gaussianity introduced due to the polarization fraction tensor transformation.
@@
The first set of maps contains isotropic small-scale structure. We construct the small-scale structure by generating a Gaussian random field with power-law power spectra in $TT$, EEEE, and $BB$, using power laws fit to the power spectrum of the d10 maps on the GAL097 mask in the multipole range $[800,2000]$. This Gaussian random field is then high-pass filtered to remove power below $\ell_{cut}=200$ in both total intensity and polarization, and co-added to the large-scale dust template.
@@
For the second set of maps, we introduce a modulation in $I$, $Q$, and $U$ that mimics the effective modulation applied to $i$, $q$, and $u$ in the construction of the d10 maps. We generate modulation maps $m_{I}$ and $m_{P}$ from $IQU$, following the same procedure as done in iquiqu space (Equations 89). We multiply the Gaussian isotropic small-scale map described above with these modulation maps, and then co-add them with the large-scale dust template. However, we want to ensure that the power spectra of modulated Gaussian small scales computed on the individual sky-fraction masks are as close as possible to the d10 map. To achieve this, we adjust the modulation maps $m_{I}$ and $m_{P}$ by applying different multiplicative factors to non-overlapping regions of the sky. We multiply $m_{I}$ and $m_{P}$ on the GAL40 mask by scalar factors that adjust the mean power spectrum of the modulated small-scale structure to be equal to the power spectrum of d10 on the same mask. We repeat this process for each successively larger sky-fraction mask, applying the approximately order-unity multiplication factor to the non-overlapping sky region at each iteration.
@@
We thus consider three sets of maps with different small scales co-added: model d10, a map with purely Gaussian small-scale structure, and a map with modulated Gaussian small-scale structure. We refer to these three maps as poltens_ss, Gaussian_ss, and Gaussian_mod_ss, respectively.
We apply a high-pass filter with $\ell_{min}=200$, using a smooth function similar to Equation 11, to retain only the small scales of these maps. We calculate the MFs both on the sphere and in several selected regions of the sky projected into a Cartesian projection.
V.5.1 Minkowski Functionals on the sphere
@@
Following the algorithm in Grewal:2022, we calculate the MFs for the three $Q$ maps on the sphere, i.e., in HEALPix format, on the GAL080 mask. We first normalize the maps by dividing each map by its standard deviation, and compute the MFs for iso-intensity contours in the range [3,3][-3,3] (Figure 13). The MFs of the corresponding $U$ maps look very similar to $Q$ maps and are not shown here.
@@
Figure 13 shows that when averaging over a large sky area, the MFs of Gaussian_mod_ss and poltens_ss are almost identical, while the MFs of Gaussian_ss differ substantially. The difference in MFs between poltens_ss and Gaussian_ss indicates the existence of non-Gaussianity in poltens_ss, but the similarity between poltens_ss and Gaussian_mod_ss demonstrates that the non-Gaussianity in poltens_ss comes from the anisotropy in the maps, which originates from the modulation, rather than from the polarization fraction tensor transformation.
@@
Figure 14: Zoom-in plots of an intermediate Galactic latitude patch centered at $(l,b)=(-15^{\circ}$, $45^{\circ})$. The $I$, $Q$, and $U$ maps (top to bottom) constructed using Gaussian_mod_ss and poltens_ss are shown in the first two columns, respectively. The latter is the d10 map. The two rightmost columns show these maps high-pass filtered to preserve only scales with $\ell>200$. The colorbar on the left indicates the pixel values in the left-most two columns in the units of $\mu\rm K_{RJ}$ and the colorbar on the right is for the last two columns.
@@
Figure 15: Minkowski Functionals as a function of the threshold $\rho$ for one realization of $I$, $Q$ and $U$ small scales ($\ell>200$) in the patch centered on $(l,b)=(-15^{\circ}$, $45^{\circ})$ in Galactic coordinates. Each row shows three kinds of Minkowski Functionals. The blue dashed one is for Gaussian_mod_ss while the orange solid one is from poltens_ss. We also show the Gaussian_ss in dashed gray line as a comparison.
V.5.2 Minkowski Functionals on small regions
@@
We consider an intermediate-latitude $20^{\circ}\times 20^{\circ}$ region centered at $(l,b)=(-15^{\circ}$, $45^{\circ})$ to determine whether significant differences in the MFs between Gaussian_mod_ss and poltens_ss sets of maps exist in small regions of sky. Those maps are shown in Figure 14. We can see by eye that poltens_ss contains structure that is not present in the Gaussian_mod_ss maps. We calculate the MFs of these small-scale maps, following Mantz:2008 for the calculation of MFs for a square patch. Before the calculation, we also rescale all the small scales linearly to be in the range [1,1][-1,1].
@@
Figure 15 shows the MFs of the Gaussian_mod_ss and poltens_ss maps presented in Figure 14. In contrast with the large-area results presented in Figure 13, in this case we do measure a departure of the poltens_ss MFs from the Gaussian_mod_ss ones. This means that the polarization fraction tensor transformation introduced non-Gaussian small-scale structure, distinct from pure modulation effects, that is detectable over small sky regions. We conclude that the level of induced non-Gaussianity differs from region to region and does not significantly impact the statistical properties when averaged over large sky fractions.
VI Future Outlook
@@
The Galactic emission templates employed in this work are derived from component separation of the microwave sky and thus are subject to both noise and fitting degeneracies. This is particularly evident in the spectral parameter maps. The $T_{d}$ and $\beta_{d}$ template maps have a strong anti-correlation that can arise from fitting degeneracies, particularly at low signal-to-noise ratios (Shetty:2009). These maps also exhibit correlations with the dust intensity maps, particularly in diffuse, low-signal regions. While all dust models developed in this work respect current observational constraints on frequency decorrelation (see Section V.3), it is likely that some variation in the spectral parameter maps is attributable to noise rather than true astrophysical variation. Construction of improved spectral parameter maps is feasible by employing improved algorithms on existing data (e.g., Watts:2024) and should be a high priority for future work.
@@
By imposing a uniform cut-off in $\ell$ over which data-driven templates are employed, the models developed here are discarding high signal-to-noise information at small scales over portions of the sky where emission is bright. Future methodologies would ideally retain information from the templates on as small a scale as possible, with the cut-off scale varying across the sky. Artifacts from the splicing of large-scale templates with small-scale synthetic structures, including the ratio of spatial power across the modulation scale discussed in Section V.2.2, should be further mitigated. We note that the current methodology is fundamentally not designed to be able to reproduce the observed power spectra of all quantities over all sky areas, but future approaches could improve the behavior of the power spectrum beyond the modulation scale and for smaller sky masks. Further, particularly as data are being collected over partial sky areas by sensitive ground-based experiments, methods to incorporate partial sky templates should be developed.
@@
The current models do not explicitly impose any non-zero parity-odd correlations in the polarized dust emission, i.e., the $TB$ and $EB$ correlations are zero in the power spectra that are used to extrapolate the small-scale dust emission structure. However, analysis of the Planck polarized dust emission finds a significant nonzero $TB$ correlation at $\ell\lesssim 600$ (planck2016-l11A; Weiland:2020). A proposed physical mechanism for the origin of the nonzero $TB$ signal is misalignment between dusty filaments and the projected magnetic field orientation (Huffenberger:2020; Clark:2021; Cukierman:2023): this picture also predicts the sign and amplitude of an expected nonzero $EB$ correlation. Future work could incorporate this parity-odd contribution to the polarized dust emission (e.g., Hervias-Caimapo:2025). Such models would be of particular use for the development of analysis techniques that seek to constrain signatures of cosmic birefringence or other parity-violating physics in the CMB (e.g., Choi:2020; Minami:2020; Eskilt:2022; Ade:2025a).
@@
Future work could also improve the physical realism of the small-scale emission structure. The structure of polarized dust emission on small scales is highly filamentary (e.g., Clark:2015; Halal:2024b). Realizations of small-scale structure that have the particular character of non-Gaussianity that is realized in the sky could be generated using, for example, generative adversarial neural networks (Krachmalnicoff:2021; Yao:2024), a transformer-based super-resolution approach (Halal:2024thesis), or scattering-transform-based techniques (Regaldo-SaintBlancard:2020; Delouis:2022; Mousset:2024), which could be informed by the non-Gaussian statistics of diffuse ISM gas tracers (Lei:2023; Lei:2025). Forthcoming high-angular resolution dust maps from, e.g., SO (Ade:2019), the South Pole Observatory (TheBICEP/KeckCollaboration:2024), or CCAT (CCAT-PrimeCollaboration:2023) could also be used to inform a statistical description of small-scale dust structure. Another approach is to explicitly model the observed emission as arising from a superposition of many interstellar filaments, perhaps informed by the structure of interstellar neutral hydrogen emission (Clark:2019), or as a flexible parametric model (Hervias-Caimapo:2022; Hervias-Caimapo:2025).
@@
As our knowledge of diffuse ISM physics improves, there are further opportunities for more sophisticated modeling of the Galactic emission. A recently discovered link between the mass fraction of neutral hydrogen in the cold neutral medium and the 353 GHz dust polarization fraction implies that dust polarization models could be further improved by incorporating information on the phase structure of the neutral ISM (Lei:2024). It was also recently found that the three-dimensional structure of nearby interstellar dust has a measurable imprint on the dust polarization structure at CMB-relevant frequencies (Halal:2024a). Thus, as models of the 3D dust distribution (e.g., Edenhofer:2024; Zhang:2025) improve, it may be possible to incorporate these data products into models of the dust polarization field. Starlight polarization data can further constrain the plane-of-sky magnetic field orientation of nearby dust structures (e.g., Pelgrims:2024; Panopoulou:2025). Although our focus is on data-driven sky modeling, numerical simulations (e.g., Kim:2019; Stalpes:2024) could plausibly inform modeling choices for aspects of the models that are poorly constrained by data, e.g., higher-order statistics of the dust polarization (e.g., Portillo:2018; Williamson:2024).
@@
In addition to the small-scale structure of the diffuse dust and synchrotron emission, the real sky at these frequencies contains emission from discrete sources that are not explicitly included in our models. These include cold clumps (Clancy:2023), as well as planetary nebulae, supernova remnants, pulsar wind nebulae, and so forth (Naess:2020; Guan:2021). Nearby galaxies, while not technically within the scope of Galactic emission modeling, should also be incorporated as they are not included in existing simulations of the extragalactic sky.
VII Recommended Model Suite